# magnetohydrodynamics (MHD) — turbulence — diffusion — protoplanetary disks

### v1

@@ 4.2.1 The Active Zone
In the early phase ($t\lesssim 1300~{}t_{\mathrm{K0}}$), the mass of the active zone is lost both vertically and radially with significant time variations. In the radial directions, the mass of the active zone flows out more from $R=R_{+}=R_{\mathrm{MRI}}$ than from $R=R_{-}=R_{\mathrm{buf}}$. The mass ejection rate by the disk wind $(\dot{M}_{\mathrm{ac}})_{z}$ is comparable to $(\dot{M}_{\mathrm{ac}})_{R+}$. The mass transport rate through $R=R_{+}$ is investigated in Section 4.2.2 in detail. This indicates that the vertical mass loss is as important as the radial mass transport.
@@
In the later phase ($t\gtrsim 1300~{}t_{\mathrm{K0}}$), $(\dot{M}_{\mathrm{ac}})_{R-}$ and $(\dot{M}_{\mathrm{ac}})_{R+}$ approach zero. The vertical mass flow mainly removes the mass of the active zone. Where does the mass ejected vertically from $z=\pm z_{\mathrm{atm}}(R)$ go? Since the radial velocity in $|z|>z_{\mathrm{atm}}$ above the active zone is negative (the bottom panels of Figure 23 and Figure 25), the disk wind falls onto the center. In order to take into account mass accretion at high latitudes at $R=R_{\mathrm{buf}}$, we measure the mass accretion rate $(\dot{M}_{\mathrm{ac}})_{\mathrm{buf}}$ through the sphere $r=R_{\mathrm{buf}}$, where we take into account only the regions where $\rho>10^{-2}$ to remove the funnel regions. Figure 27 shows that $(\dot{M}_{\mathrm{ac}})_{\mathrm{buf}}$ is comparable to $(\dot{M}_{\mathrm{ac}})_{z}$ in all the times for LowRes and HighRes runs. This indicates that most mass ejected vertically falls onto the center (Takasao et al., 2018).
@@ 4.2.2 The Transition Zone
The mass flux from the active zone to the transition zone $(\dot{M}_{\mathrm{tr}})_{R-}=(\dot{M}_{\mathrm{ac}})_{R+}$ shows significant quasi-periodic time variations in the early phase of $t\lesssim 1300t_{\mathrm{K0}}$; time periods with large and small $(\dot{M}_{\mathrm{tr}})_{R-}$ are periodically repeated. The rate of the mass supply passing through $R=R_{-}=R_{\mathrm{MRI}}$ is positive most of the time. This is consistent with the mass transfer rate shown in Figures 23. The time variability of $(\dot{M}_{\mathrm{tr}})_{R-}$ is attributed to the MRI activity in the active zone. For $t\gtrsim 1300t_{\mathrm{K0}}$, the quasi-periodic time variations in $(\dot{M}_{\mathrm{tr}})_{R-}$ disappear and $(\dot{M}_{\mathrm{tr}})_{R-}$ gradually decreases with time. This corresponds to the decrease in the MRI activity in the active zone shown in Figure 37.
Figure 26b shows that the mass transfer rate $(\dot{M}_{\mathrm{tr}})_{R+}$ is positive and does not show a significant time variation. This is consistent with the fact that the gas is transferred from the dead zone to the transition zone by the quasi-steady surface accretion flow.
Figure 28: Vertical mass outflow rate from $R$ to $R+dR$ as a function of radius. Panel (a) compares the results for LowRes and HighRes, and they are averaged over $400\leq t/t_{\mathrm{K0}}\leq 500$. Panel (b) displays the time evolution of $d\dot{M}_{\mathrm{wind}}/d\ln R$ by showing the result averaged over three different time intervals, $500\leq t/t_{\mathrm{K0}}\leq 1000$, $1000\leq t/t_{\mathrm{K0}}\leq 1500$, and $1500\leq t/t_{\mathrm{K0}}\leq 2000$. The gas accretion rate in the coherent zone is shown by the magenta line. All the quantities are shown in the code units.
Figure 26b shows that the vertical mass supply $(\dot{M}_{\mathrm{tr}})_{z}$ stays almost zero. This does not mean that there are no mass transfer at each radius. To investigate the radial profile of the mass ejection rate by the disk wind, in Figure 28 we plot the mass ejection rate from the surfaces of $z=\pm z_{\mathrm{atm}}$ from $R$ to $R+dR$ that is given by
@@ 4.2.3 The Coherent Zone
Unlike the active and transition zones, the change rate of the coherent zone mass ($R_{\eta}\leq R\leq R_{\mathrm{ch}}$) appears to approach zero in the long-term evolution (Figure 26c). The mass supply through $R=R_{+}=R_{\mathrm{ch}}$ almost balances with the mass loss through $R=R_{-}=R_{\eta}$ and disk wind. We note that our calculation time may not be long enough for $(\dot{M}_{\mathrm{co}})_{R+}$ to reach a quasi-steady state. About half of the mass supplied through $R=R_{+}=R_{\mathrm{ch}}$ is ejected by the disk wind, and the remaining half is transferred to the transition zone since $(\dot{M}_{\mathrm{co}})_{z}\sim(\dot{M}_{\mathrm{co}})_{R-}$ (Figure 26c).
@@
which indicates a mass ejection efficiency because $\dot{M}_{\mathrm{wind}}$ obtained from the radial integration of $d\dot{M}_{\mathrm{wind}}/d\ln R$ increases more rapidly compared with the accretion rate $\dot{M}^{z_{\mathrm{atm}}}$ as the radial extent considered in the integration for larger $\xi$.
@@ 4.3.1 The Active Zone
$\left(\langle({E}_{\phi})_{\mathrm{I}}\rangle_{H}\right)_{\mathrm{turb}}=\langle(v_{R}-\langle v_{R}\rangle_{H,v})\delta B_{z}\rangle_{H}-\langle(v_{z}-\langle v_{z}\rangle_{H,v})\delta B_{R}\rangle_{H},$ (68)
@@ 6 Summary
Our simulations have demonstrated that the complicated structures formed by the non-ideal MHD effects affect the mass, angular momentum, and magnetic flux transport as well as dust accumulation. We are planning to perform wider parameter search to investigate such structures systematically. In addition, we will include the Hall effect in our future work.
@@
Numerical computations were carried out on supercomputer Fugaku provided by the RIKEN Center for Computational Science (Project ids: hp210164, hp220173), Oakforest-PACS at Joint Center for Advanced High Performance Computing (Project ids: hp190088, hp200046, hp210113), and Cray XC50 at the CfCA of the National Astronomical Observatory of Japan. This work was supported in part by the Ministry of Education, Culture, Sports, Science and Technology (MEXT), Grants-in-Aid for Scientific Research, JP19K03929 (K.I.), JP16H05998 (K.T. and K.I.), JP21H04487, JP22KK00443 (K.T., K.I., and S.T.), JP23H01227 (S.O.), JP17H01105, JP21H00033, and JP22H01263 (T.K.S.). This research was also supported by MEXT as “Exploratory Challenge on Post-K computer” (Elucidation of the Birth of Exoplanets [Second Earth] and the Environmental Variations of Planets in the Solar System) and “Program for Promoting Researches on the Supercomputer Fugaku” (Toward a unified view of the universe: from large scale structures to planets, JPMXP1020200109).
@@ Appendix A Resolution Dependence of the $\alpha$ Parameter
Figure 36 compares the time evolution of $\langle\alpha_{R\phi}\rangle$ (Equation (36) between LowRes and HighRes runs. Both models show that the MRI turbulence is saturated after $t\sim 20t_{\mathrm{K}}(R=0.3~{}\mathrm{au})$. The saturation level of $\langle\alpha_{R\phi}\rangle_{H}$ is about twice larger for HighRes than for LowRes runs, consistent with the fact that LowRes run does not resolve the MRI turbulence sufficiently. The saturation level $\langle\alpha_{R\phi}\rangle_{H}\sim 2.8\times 10^{-2}$ for HighRes run is comparable to those found in local shearing-box simulations (e.g., Hawley et al., 1995; Sano et al., 2004). We confirmed that HighRes run is expected to provide converged results by measuring the so-called quality factors (Noble et al., 2010).

### v2

@@ 4.2.1 The Active Zone
Early Evolution ($t\lesssim 1000~{}t_{\mathrm{K0}}$)
In the early phase ($t\lesssim 1000~{}t_{\mathrm{K0}}$), the mass of the active zone is lost both vertically and radially with significant time variations. The fact that the mass of the active zone is lost through $R=R_{\mathrm{MRI}}$ ($(\dot{M}_{\mathrm{ac}})_{R+}<0$) is consistent with the positive mass transfer rate at $R=R_{\mathrm{MRI}}$ in Figures 23c and 23d.
In the radial directions, the mass of the active zone flows out more from $R=R_{+}=R_{\mathrm{MRI}}$ than from $R=R_{-}=R_{\mathrm{buf}}$. The mass ejection rate by the disk wind $(\dot{M}_{\mathrm{ac}})_{z}$ is comparable to $(\dot{M}_{\mathrm{ac}})_{R+}$. This indicates that the vertical mass loss is as important as the radial mass transport.
Why does $(\dot{M}_{\mathrm{ac}})_{R+}$ exhibit quasi-periodic variation even though the MRI disappears around $R=R_{\mathrm{MRI}}$ (Figure 24)? The quasi-periodic variations generated in the atmospheres of the active zone disturb the magnetic field around $R=R_{\mathrm{MRI}}$. Figure 26b shows that the time variations of $\langle B_{\phi}\rangle$ at $z=2H$ at two different radii. As a representative radius in the active zone, $R=0.3~{}\mathrm{au}$ is chosen because the quasi-periodic variations caused by the MRI are seen in $0.2~{}\mathrm{au}\lesssim R\lesssim 0.4~{}\mathrm{au}$. At $R=0.3~{}\mathrm{au}$, $\langle B_{\phi}\rangle$ at $z=2H$ is positive for a longer time than it is negative because it is amplified from negative $B_{R}$ that is generated by dragging of the magnetic fields toward the central star (Suzuki & Inutsuka, 2014; Takasao et al., 2018). The amplified positive $\langle B_{\phi}\rangle$ suddenly breaks and the sign of $\langle B_{\phi}\rangle$ is flipped. Afterwards, the positive $\langle B_{\phi}\rangle$ is amplified again. This cycle is repeated in a quasi-periodic manner in the active zone (Figure 37). When the sign of $\langle B_{\phi}\rangle$ reverses and reverts in the active-zone atmospheres, the disturbances propagate outwards and affect the magnetic fields around $R=R_{\mathrm{MRI}}$. This can be seen in the fact that $\langle B_{\phi}\rangle(R=R_{\mathrm{MRI}},z=2H)$ suddenly drops slightly after $\langle B_{\phi}\rangle(R=0.3~{}\mathrm{au},z=2H)$ becomes negative in Figure 26b. The quasi-periodic variations in the magnetic fields at $R=R_{\mathrm{MRI}}$ produce the fluctuations of $(\dot{M}_{\mathrm{ac}})_{R+}$.
Comparison between Figures 26a and 26b indicates that the time variation of the vertical mass transfer rate $(\dot{M}_{\mathrm{ac}})_{z}$ is also correlated with that of $\langle B_{\phi}\rangle$ at $R=0.3~{}\mathrm{au}$. This is because the gas is ejected radially obliquely upward when the sign of $\langle B_{\phi}\rangle$ returns to positive in the northern hemisphere.
The variations of $(\dot{M}_{\mathrm{ac}})_{R-}$ are not quasi-periodic and correlate little with $(\dot{M}_{\mathrm{ac}})_{R+}$ and $(\dot{M}_{\mathrm{ac}})_{R+}$. The variations of $(\dot{M}_{\mathrm{ac}})_{R-}$ are also likely to be produced by the MRI activity, but it may be affected by the buffer region and the inner boundary conditions (Section 2.5).
Later Evolution ($t\gtrsim 1000~{}t_{\mathrm{K0}}$)
The time evolution of $(\dot{M}_{\mathrm{ac}})_{R+}$ transitions to a stationary state because the magnetic field around $R=R_{\mathrm{MRI}}$ is no longer subject to quasi-periodic disturbances from the atmospheres of the active zone (Figure 26b). As mentioned in Section 4.1.2, $(\dot{M}_{\mathrm{ac}})_{R+}$ approaches zero because the mid-plane gas accretion is almost compensated by the outflow in higher latitudes.
@@
The vertical mass flow $(\dot{M}_{\mathrm{ac}})_{z}$ mainly determines the time evolution of the mass of the active zone because $(\dot{M}_{\mathrm{ac}})_{R-}$ and $(\dot{M}_{\mathrm{ac}})_{R+}$ approach zero (Figure 26a). Where does the mass ejected vertically from $z=\pm z_{\mathrm{atm}}(R)$ go? Since the radial velocity in $|z|>z_{\mathrm{atm}}$ above the active zone is negative (the bottom panels of Figure 23 and Figure 25), the disk wind falls onto the center. In order to take into account mass accretion at high latitudes at $R=R_{\mathrm{buf}}$, we measure the mass transfer rate $(\dot{M}_{\mathrm{ac}})_{\mathrm{buf}}$ through the sphere $r=R_{\mathrm{buf}}$, where we take into account only the regions where $\rho>10^{-2}$ to remove the funnel regions. Figure 27 shows that $(\dot{M}_{\mathrm{ac}})_{\mathrm{buf}}$ is comparable to $(\dot{M}_{\mathrm{ac}})_{z}$ in all the times for LowRes and HighRes runs. This indicates that most mass ejected vertically falls onto the center (Takasao et al., 2018).
@@ 4.2.2 The Transition Zone
Figure 26c shows that the mass transfer rate $(\dot{M}_{\mathrm{tr}})_{R+}$ is positive and does not show a significant time variation. This is consistent with the fact that the gas is transferred from the coherent zone to the transition zone by the quasi-steady surface accretion flow.
Figure 28: Vertical mass outflow rate from $R$ to $R+dR$ as a function of radius. Panel (a) compares the results for LowRes and HighRes, and they are averaged over $400\leq t/t_{\mathrm{K0}}\leq 500$. Panel (b) displays the time evolution of $d\dot{M}_{\mathrm{wind}}/d\ln R$ by showing the result averaged over three different time intervals, $500\leq t/t_{\mathrm{K0}}\leq 1000$, $1000\leq t/t_{\mathrm{K0}}\leq 1500$, and $1500\leq t/t_{\mathrm{K0}}\leq 2000$. The mass transfer rate in the coherent zone is shown by the magenta line. All the quantities are shown in the code units.
Figure 26c shows that the vertical mass supply $(\dot{M}_{\mathrm{tr}})_{z}$ stays almost zero. This does not mean that there are no mass transfer at each radius. To investigate the radial profile of the mass ejection rate by the disk wind, in Figure 28 we plot the mass ejection rate from the surfaces of $z=\pm z_{\mathrm{atm}}$ from $R$ to $R+dR$ that is given by
@@ 4.2.3 The Coherent Zone
Unlike the active and transition zones, the change rate of the coherent zone mass ($R_{\eta}\leq R\leq R_{\mathrm{ch}}$) appears to approach zero in the long-term evolution (Figure 26d). The mass supply through $R=R_{+}=R_{\mathrm{ch}}$ almost balances with the mass loss through $R=R_{-}=R_{\eta}$ and disk wind. We note that our calculation time may not be long enough for $(\dot{M}_{\mathrm{co}})_{R+}$ to reach a quasi-steady state. About half of the mass supplied through $R=R_{+}=R_{\mathrm{ch}}$ is ejected by the disk wind, and the remaining half is transferred to the transition zone since $(\dot{M}_{\mathrm{co}})_{z}\sim(\dot{M}_{\mathrm{co}})_{R-}$ (Figure 26d).
@@
which indicates a mass ejection efficiency because $\dot{M}_{\mathrm{wind}}$ obtained from the radial integration of $d\dot{M}_{\mathrm{wind}}/d\ln R$ increases more rapidly compared with the mass transfer rate $\dot{M}^{z_{\mathrm{atm}}}$ as the radial extent considered in the integration for larger $\xi$.
@@ 4.3.1 The Active Zone
In this section, we will show that the magnetic flux drifts inward (outward) in inner (outer) regions of the active zone in the early phase evolution. In the later phase, the drift velocities oscillate between positive and negative with time in the inner active zone while the magnetic flux drifts inward in the outer active zone.
@@
$\left(\langle({E}_{\phi})_{\mathrm{I}}\rangle_{H}\right)_{\mathrm{turb}}$ $=$ $\langle(v_{R}-\langle v_{R}\rangle_{H,v})\delta B_{z}\rangle_{H}$ (68)
$-\langle(v_{z}-\langle v_{z}\rangle_{H,v})\delta B_{R}\rangle_{H},$
@@ 5.1.2 The Transition and Coherent Zones
5.1.3 Vertical Shear Instability
Since the locally isothermal equation of state is used in our simulations (Section 2.1), the VSI may occur (e.g., Urpin & Brandenburg, 1998; Nelson et al., 2013). In this section, we will show that our simulation time is too short to see non-linear growth of the VSI in the transition and coherent zones.
Vertically global linear analyses show that the VSI modes can be divided into the two body modes (corrugation and breathing modes) that have large amplitudes near the mid-plane and the surface modes that emerge when the boundaries are set in the vertical direction (Nelson et al., 2013; Barker & Latter, 2015; Umurhan et al., 2016). In the nonlinear stage, the fundamental corrugation mode dominates over the other modes (e.g., Nelson et al., 2013).
In all the regions, the VSI does not appear to grow in our simulations. In the active zone, the VSI is suppressed by magnetic fields (Latter & Papaloizou, 2018; Cui & Lin, 2021; Latter & Kunz, 2022). Also in the transition and coherent zones, our simulations do not show the non-linear development of the corrugation modes.
[…]
The surface modes, which develops near the vertical boundaries, is suppressed due to the limited vertical extent ($|z|\leq H$) when the radial wavelength is less than $\sim 40(\epsilon^{2}R)^{-1}$ in the reduced model. Although the surface modes would emerge for higher $k_{\mathrm{R}}$ at higher growth rates, the radial resolution of our simulations is not sufficient to resolve them.
@@ 5.4 Implications for Protoplanetary Disks Around Solar-type Stars
Jankovic et al. (2021, 2022) investigated steady-state axisymmetric disk structures considering detailed physical and chemical processes. They found that the temperature structure is more complex than our adopted one (Equation (10)). This will affect the shape of the dead zone boundaries. They derived a similar $R_{\mathrm{MRI}}$ as Equation (71) even when they took into account the disk structures modified by convective instability.
@@ 6 Summary
Our simulations have demonstrated that the complicated structures formed by the non-ideal MHD effects affect the mass, angular momentum, and magnetic flux transport as well as dust accumulation. We are planning to perform wider parameter search to investigate such structures systematically.
There are two main caveats in our study. One is that the Hall effect is neglected for simplicity. In our model setup, the diffusion coefficient of the Hall effect is as large as $\eta_{\mathrm{O}}$ and $\eta_{\mathrm{A}}$ near the dead-zone inner boundaries. We will include the Hall effect because it can change the disk structure significantly (Bai & Stone, 2017; Béthune et al., 2017; Martel & Lesur, 2022). The other is that we use the locally isothermal equation of state with a simple temperature profile (Equation (10)). The temperature profile is more complicated near the dead-zone inner boundaries when the radiative transfer is considered (e.g., Flock et al., 2017; Gressel et al., 2020). This can change the spatial distributions of the diffusion coefficients and the thermal evolution of the gas.
@@
Numerical computations were carried out on supercomputer Fugaku provided by the RIKEN Center for Computational Science (Project ids: hp210164, hp220173), Oakforest-PACS at Joint Center for Advanced High Performance Computing (Project ids: hp190088, hp200046, hp210113), and Cray XC50 at the CfCA of the National Astronomical Observatory of Japan.
Funding
This work was supported in part by the Ministry of Education, Culture, Sports, Science and Technology (MEXT), Grants-in-Aid for Scientific Research, JP19K03929 (K.I.), JP16H05998 (K.T. and K.I.), JP21H04487, JP22KK0043 (K.T., K.I., and S.T.), JP22K14074 (S.T.), JP20H00182, JP23H00143, JP23K25923, JP23H01227 (S.O.), JP17H01105, JP21H00033, and JP22H01263 (T.K.S.). This research was also supported by MEXT as “Exploratory Challenge on Post-K computer” (Elucidation of the Birth of Exoplanets [Second Earth] and the Environmental Variations of Planets in the Solar System) and “Program for Promoting Researches on the Supercomputer Fugaku” (Toward a unified view of the universe: from large scale structures to planets, JPMXP1020200109).
@@ Appendix A Resolution Dependence of the $\alpha$ Parameter
Figure 36 compares the time evolution of $\langle\alpha_{R\phi}\rangle_{H}$ (Equation (36) between LowRes and HighRes runs. Both models show that the MRI turbulence is saturated after $t\sim 20t_{\mathrm{K}}(R=0.3~{}\mathrm{au})$. The saturation level of $\langle\alpha_{R\phi}\rangle_{H}$ is about twice larger for HighRes than for LowRes runs, consistent with the fact that LowRes run does not resolve the MRI turbulence sufficiently. The saturation level $\langle\alpha_{R\phi}\rangle_{H}\sim 2.8\times 10^{-2}$ for HighRes run is comparable to those found in local shearing-box simulations (e.g., Hawley et al., 1995; Sano et al., 2004). We confirmed that HighRes run is expected to provide converged results by measuring the so-called quality factors (Noble et al., 2010).